# Fully Distributed Fog Load Balancing with Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Real-time Internet of Things (IoT) applications require real-time support to handle the ever-growing demand for computing resources to process IoT workloads. Fog Computing provides high availability of such resources in a distributed manner. However, these resources must be efficiently managed to distribute unpredictable traffic demands among heterogeneous Fog resources. This paper proposes a fully distributed load-balancing solution with Multi-Agent Reinforcement Learning (MARL) that intelligently distributes IoT workloads to optimize the waiting time while providing fair resource utilization in the Fog network. These agents use transfer learning for life-long self-adaptation to dynamic changes in the environment. By leveraging distributed decision-making, MARL agents effectively minimize the waiting time compared to a single centralized agent solution and other baselines, enhancing end-to-end execution delay. Besides performance gain, a fully distributed solution allows for a global-scale implementation where agents can work independently in small collaboration regions, leveraging nearby local resources. Furthermore, we analyze the impact of a realistic frequency to observe the state of the environment, unlike the unrealistic common assumption in the literature of having observations readily available in real-time for every required action. The findings highlight the trade-off between realism and performance using an interval-based Gossip-based multi-casting protocol against assuming real-time observation availability for every generated workload.
@@
Internet of Things, Fog Computing, Load Balancing, Reinforcement Learning, Multi-Agent Learning
@@ I Introduction
Load balancing in Fog computing environments plays a crucial role in ensuring optimal resource utilization and timely processing of IoT data; this allows for minimizing task completion time and improving overall network performance. Traditional load-balancing techniques often struggle with the dynamic nature of Fog environments, characterized by heterogeneous resources, limited communication bandwidth, and fluctuating workloads. This is where Reinforcement Learning (RL) comes into play, as it allows agents to learn optimal decision-making policies through interaction with the environment. However, managing Fog and IoT resources must be done at the network edge, in a distributed manner. This will avoid the need for distant decisions, taking full advantage of the distributed computing capabilities of Fog systems [1].
Unlike centralized RL-based load-balancing solutions, fully distributed Multi-Agent RL (MARL) overcomes scalability limitations by allowing autonomous agents to learn from their local observations and interactions. These agents can be deployed in network Access Points (APs), i.e., Base Stations, to distribute the load their associated end devices generate. This paper presents a novel fully distributed approach for Fog load balancing using MARL. Our approach introduces distributed self-learning agents capable of efficiently distributing IoT load across the Fog environment. These agents incorporate self-adaptation mechanisms, leveraging practical lifelong transfer learning to adapt to dynamic environmental changes seamlessly. This distributed solution outperforms the centralized lifelong learning agent proposed in our previous work [2].
Available MARL solutions in the literature provide collaboration between RL agents through one or more centralized components, like a centralized policy and a centralized experience replay buffer, which is not practical for a global-scale Fog load-balancing solution. Without fully decentralizing decision-making, the communication overhead grows exponentially with the number of agents and Fog nodes in the system, especially when the agents need to communicate with each other. In addition, proposing fully independent agents lowers the state and action spaces of individual agents to reduce the computational complexity of the solution. This is done by considering smaller sets of Fog nodes, e.g., within a certain proximity, as possible candidates for each group of agents (see Fig. 1).
@@
Fig. 1 shows the formation of smaller collaboration regions, where the group of agents accessing the same set of candidate Fog nodes can form a collaboration region. Each region contains one or more agents that need to distribute the load between one or more candidate Fog nodes. In this setup, an agent can only be part of a single collaboration region. Candidate Fog nodes can be shared between two or more overlapping regions, while an isolated collaboration region does not share its set of candidate Fog nodes with other regions. Allowing the agents to independently, i.e., blindly, learn to collaborate in each collaboration region leads to the global optimization of the whole system.
The independent agents must learn to optimize the same objective of minimizing the overall system waiting delay. To avoid unnecessary communication overhead, each agent observes the same state from the environment while independently taking actions and receiving rewards for those actions. Hence, our independent distributed agents can work without sharing their knowledge, experience, or actions, avoiding the need for a centralized coordination entity and inter-agent communication. In addition, a realistic multi-cast approach for observing the environment is proposed in this work, compared to unrealistic real-time observations often considered in the literature.
@@
This paper proposes the first fully distributed, efficient, and self-adapting MARL load distribution solution for realistic global-scale Fog Computing implementations without centralized components or inter-agent communication. Through empirical evaluation, we show the effectiveness of our novel truly distributed MARL-based load-balancing solution and compare it with baseline solutions and a centralized state-of-the-art RL solution. We also show the performance of observing the environment using realistic multi-cast intervals compared to unrealistic real-time observations. Hence, the key contributions of this paper can be summarized as follows:
  • Propose fully distributed self-learning LB RL agents that efficiently distribute IoT load into a Fog Computing environment.
  • Augment the independent distributed agents with lifelong self-adaptation to dynamic environment changes using transfer learning.
  • Provide a global-scale solution by allowing the agents to work independently in small collaboration regions, unlike centralized architectures.
  • Evaluate realistic interval-based collection of observations using Gossip multi-casting protocol compared to the common practice in the literature of unrealistically providing this information in real-time.
@@ II Related Work
Jain et al. [3] proposed training an agent for each Fog node to mitigate the delay of decision-making in centralized solutions. However, they trained all the agents offline on a centralized cloud data center to later distribute the trained models to each Fog node to make offloading decisions locally with minimal delay. However, training these agents using this approach, and retraining them in case of changes in the environment, requires the collection of local experience trajectories of each agent, i.e., observations, actions, rewards, and next observations. Transferring such information to a centralized entity to retrain these agents is inefficient compared to local retraining, especially for highly dynamic environments that keep changing over time.
@@
Determining the optimal timing for retraining these agents in the centralized entity presents a critical trade-off between decision accuracy and the significant retraining overhead. In addition, an RL decision can only be made when the observation of the current step is available, along with the reward of the previous step. Observing the state, for each decision, of every Fog node in large-scale Fog networks will cause a significant delay, especially with frequently generated workloads [4]. This delay is often not considered in the literature when proposing load-balancing solutions with real-time decision-making. Hence, the effectiveness of these solutions in practice can not be guaranteed without considering the time needed to transfer Fog load provisioning information to all RL agents.
@@
Gao et al. [5] used MARL to address the scalability problem of RL-based load-balancing solutions. They provide task offloading and resource allocation in large-scale heterogeneous Multi-access Edge Computing (MEC) systems. In their design, each edge server is regarded as an actor agent while still considering a centralized critic, where the critic is a neural network used to evaluate the actions of actor networks [6]. This centralized critic facilitates the coordination of offloading policies among those distributed actors. Besides not being fully distributed, the authors assumed that observations are readily available when the agents need to make assignment decisions. This is a common unrealistic assumption in the literature, where the delay of exchanging Fog status information to every agent is often ignored.
@@
Baek et al. [7] proposed a centralized critic network with distributed individual actor networks. Those distributed agents are located in APs and they cooperatively learn to solve the joint link and server load-balancing problem. The policy of the distributed cooperative agents is optimized using parameter-sharing and distributed execution; the agents communicate among themselves while evaluating a common value function. The authors assume that updated policy parameters are shared between the agents by exchanging communication messages over limited bandwidth channels. However, they did not show the transmission delay associated with those communication messages, which augments when scaling up the solution for a larger-scale Fog environment. Such a delay can have a significant impact on the collaboration of the agents due to sharing outdated parameters.
@@
Similarly, Goudarzi et al. [8] proposed a MARL framework to improve the execution delay of heterogeneous Fog services using multiple actors and a centralized critic. They deployed distributed agents across different Fog nodes to work in parallel, i.e., brokers that make service offloading decisions. Then, they train a centralized service offloading model, i.e., the learner of the target policy, using the experience trajectories generated by the distributed actors. But again, they did not consider the delay caused by exchanging such information in a large-scale environment with high request rates. In addition, they did not consider the delay when updating the common value function of each actor from the centralized critic network.
@@
Wei et al. [9] proposed a scheme to improve vehicular Fog resource utilization in large-scale Vehicular Fog Computing (VFC) networks. They used MARL to decide on multi-tier task offloading in a distributed manner. Like most MARL-based solutions in the literature, they used the approach of centralized training with decentralized execution (CTDE). They used the attention mechanism to share the agents’ parameters with a joint critic network. This joint network incorporates the learning process of distributed gated actors that exploit the local computational potentials.
@@
Zhang et al. [10] also used CTDE-based MARL to reduce the long-term average delay and energy consumption while improving the throughput of dynamic Edge networks using online Request Scheduling. Centralized training was adopted to improve the long-term system performance by realizing the implicit cooperation among edge nodes. This cooperation was obtained by transmitting the replay buffer data of each agent into a shared training server to train a common policy. During distributed execution, on the other hand, only local environment information is needed for each edge node to make the request scheduling decision. However, the authors did not discuss the scope of the local information, and the delay associated with transmitting such information over the network. In a large-scale network, this delay is not tolerable for real-time decisions required by real-time applications.
@@
Lu et al. [11] proposed MARL-based dynamic scheduling to handle the massive number of service requests by the ever-increasing number of vehicles in Internet-of-Vehicle (IoV) networks. They leveraged CTDE to learn high-quality scheduling rules for the service function scheduling problem. In their architecture, the generated scheduling actions of the distributed policies are evaluated by separate critic networks deployed in a centralized controller node. Similarly, Cui et al. [12] proposed offline CTDE to cooperatively optimize the offloading strategy for computation-intensive and delay-sensitive IoV applications. Offline CTDE helped reduce the communication between the agents while increasing the stability of the environment. However, offline execution does not allow the agents to dynamically adapt to environment changes. In addition, their solution will not scale to larger networks as all their agents must observe the global state of the environment, which are assumed, unrealistically, readily available for every decision to be made.
CTDE was also used by Gao et al. [13] to predict the queuing delay of edge servers by training a centralized critic in the Cloud. To evaluate each agent’s policy, the critic network needs to collect observations and their corresponding actions from all agents in the system. Then, the agents independently make their decisions using their policies distributively. Likewise, Suzuki et al. [14] used CTDE to send learning samples from each agent to a centralized training entity to avoid the non-stationary problem of independently trained agents. A non-stationary environment is when the new state depends on the joint action of multiple agents instead of a single agent’s behavior. To learn and make offloading decisions, the agents in these approaches observe task information, network link utilization, and Fog resources utilization, which are all unrealistically assumed available in real-time for all agents decisions.
@@
Mangiaracina et al. [15] proposed adaptive MARL to manage data placement among Fog nodes considering Quality of Service (QoS) requirements. Unlike CTDE, they enabled cooperation between multiple agents using the concept of a shared repository accessible by all agents. This repository contains knowledge bases and event logs shared between those agents to improve their decision-making. The authors claim that this centralized repository does not affect the overall performance as a bottleneck in the architecture. However, they did not consider the time each agent needs to access information from this repository.
@@
To coordinate the decisions of independent agents, Yang et al. [16] proposed that the decisions of each agent must depend on shared global state information. They proposed a centralized migration scheduling model to collect server load information and calculate the migration likelihood. They estimate the global information in a centralized model to reduce the communication overhead caused by broadcasting this information from each agent to the centralized critic network. They showed a 20% gap in performance between their approach and the optimal solution but with only half the decision-making time, where the optimal solution is a centralized decision-making method having the states and actions available for all agents.
@@
Most existing MARL solutions adopt a centralized training approach, no fine-tuning for dynamic adaptation, and unrealistic assumptions for observation availability (see Table I). The centralized component in these solutions raises several challenges, especially in large-scale environments, including computational complexity, exploration of larger state and action spaces, and the communication overhead of inter-agent exchanges. Despite the efforts to mitigate these challenges in CTDE-based solutions, a truly distributed solution remains preferable to enhance scalability and decision-making time. With true distribution of decision-making, each agent makes decisions independently, without having access to the state of other collaborating agents in the system.
@@
This paper proposes a fully distributed solution that efficiently coordinates decision-making among independent agents without reliance on centralized components or inter-agent communication. This truly distributed MARL framework effectively handles the complexities of Fog environments, scales well to a global-scale solution, and minimizes communication overhead and decision latency. In addition, this work presents the first attempt to evaluate a realistic way to observe the environment and the reward of the actions of each agent based on a multi-cast Gossip protocol.

### v2

@@ Abstract
Real-time Internet of Things (IoT) applications require real-time support to handle the ever-growing demand for computing resources to process IoT workloads. Fog Computing provides high availability of such resources closer to IoT devices in a distributed manner. However, these resources must be efficiently managed to distribute unpredictable IoT traffic demands among heterogeneous Fog resources. To fully leverage these distributed Fog resources, their management must be done in a fully distributed manner, i.e., without any source of centralization management. Hence, this paper proposes a fully distributed load-balancing solution with Multi-Agent Reinforcement Learning (MARL) that intelligently distributes IoT workloads to minimize waiting delay while providing fair resource utilization in the Fog network, enhancing end-to-end execution delay. By leveraging distributed decision-making without centralized or inter-agent coordination, MARL agents can provide faster and better load distribution decisions with minimal overhead compared to a single centralized agent solution. Besides performance gain, a fully distributed solution allows for scalable Fog load balancing, where agents manage subsets of a global Fog network in smaller collaboration regions. To ensure global optimality, the agents learn to optimize a common objective, i.e., aiming to reduce workload accumulation in every Fog node. Reducing state and action spaces for each agent to a few Fog nodes in proximity makes learning easier and convergence faster. Additionally, learning independent policies by avoiding coordination encourages the agents to leverage Fog resources in proximity before seeking remote resources. In this paper, we also evaluate the impact of a realistic frequency for observing the environment instead of the common assumption of having observations readily available for every action, highlighting the trade-off between realism and performance and the solution’s applicability to real-world deployment constraints.
@@
Internet of Things, Fog Computing, Load Balancing, Reinforcement Learning, Multi-Agent Learning.
@@ I Introduction
Load balancing in Fog computing environments plays a crucial role in ensuring optimal resource utilization and timely processing of IoT data; this allows for minimizing task completion time and improving overall network performance. Traditional load-balancing techniques often struggle with the dynamic nature of Fog environments, characterized by limited heterogeneous resources and fluctuating workload demands. Reinforcement Learning (RL) can be used to learn optimal load distribution policies by interacting with such dynamic environments. To take full advantage of the distributed computing capabilities of Fog systems, load distribution decisions must be provided in a distributed manner closer to IoT devices at the network edge [1].
Fully distributed Multi-Agent RL (MARL) overcomes the scalability limitations of centralized solutions by allowing autonomous agents to independently manage smaller sets of Fog nodes in proximity instead of managing a larger network by a single agent. These gents incorporate self-adaptation by leveraging the practical lifelong transfer learning framework proposed in our previous work [2], allowing them to adapt to dynamic environmental changes seamlessly. Each agent is deployed in an IoT Access Point (AP), i.e., Base Station, to learn its own load distribution strategy for its associated IoT devices considering only the candidate Fog nodes in its region. Managing the same set of candidate Fog nodes by multiple agents forms a collaboration region (see Fig. 1), where the goal is to optimize these shared resources independently by each agent. These collaboration regions can be isolated, but they can also be overlapped by sharing candidate Fog nodes between collaboration regions as in Fig. 1.
@@
Considering small collaboration regions lowers the state and action spaces of individual agents, reducing the computational complexity of the problem and making the learning task easier for each agent. To simplify the problem further, we allow the agents to learn independent policies without collaboration, where they learn to take independent actions that fit their unique load requirements and the Fog resources closer to each agent. Learning separate load distribution strategies can be faster than learning a shared strategy for all agents, but requires a way to ensure global optimality while avoiding making the environment non-stationary for each agent. To do this in both isolated and overlapped regions, all agents learn to optimize a common objective through a common reward function, designed to reduce system-wide workload accumulation by learning to select less loaded Fog nodes.
To avoid learning a common load distribution strategy, agents must learn in a fully distributed manner without sharing any information with other agents in the system, either through a centralized coordination entity or inter-agent communication. This is because coordination eventually leads to learning a common policy, either by explicitly learning a shared policy with distributed execution of that policy or by implicitly learning a common policy through parameter and experience sharing. Parameter sharing is done by exchanging the learned parameters of the neural network (knowledge) between agents. Experience sharing can be done by exchanging the actions or the full interactions in the replay buffer.
Besides leading to a common policy, coordination causes significant communication overhead, especially with inter-agent communication with high information-sharing frequency. Most MARL solutions in the literature consider centralized collaboration through one or more centralized coordinators, reproducing the scalability problem we are trying to solve in the first place. Moreover, the overhead of observing Fog information from across the network using multiple agents is never considered in the literature. Researchers commonly assume this information is readily available in real-time for every agent’s action, which is impractical in network environments. Although this unrealistic assumption leads to theoretically optimal results, we evaluate in this work a more realistic way to observe Fog information using interval-based observations. We consider observation intervals of existing multi-casting protocols to evaluate a more realistic, although sub-optimal, performance of our solution.
@@
This paper presents a novel fully distributed, efficient, and self-adapting MARL load distribution solution for realistic global-scale Fog Computing implementations without centralized or inter-agent coordination. Through empirical evaluation, we show the effectiveness of our distributed MARL-based solution compared to the centralized RL solution proposed in our previous work [2]. We also show the performance of observing the environment using realistic multi-cast intervals of 3 seconds compared to using unrealistic real-time observations. Hence, the key contributions of this paper can be summarized as follows:
  • Propose a scalable Fog load balancing solution for a global-scale implementation using fully distributed RL agents that blindly collaborate without coordination by learning to optimize a common objective through independent actions.
  • Problem complexity is reduced by grouping the agents into smaller collaboration regions, with isolated or overlapped Fog nodes between the regions, to observe and manage smaller subsets of Fog nodes from the global network.
  • These distributed agents reduce training time by incorporating a lifelong self-adaptation framework using transfer learning, allowing the agents to efficiently distribute dynamically changing IoT load across the Fog environment.
  • Evaluate the realism-performance trade-off based on the frequency of observing Fog information from across the network, where we compare the common practice in the literature of unrealistically assuming Fog information is readily available in real-time for every action against the more realistic approach of collecting Fog information using existing interval-based multi-casting protocol, like Gossip.
@@ II Related Work
Jain et al. [3] proposed training an agent for each Fog node to mitigate the delay of decision-making in centralized solutions. However, they trained all the agents offline on a centralized cloud data center to later distribute the trained models to each Fog node, making offloading decisions locally with minimal delay. Centralized training provides the same decision-making policy for all agents despite their varying requirements, conditions, and nearby resource availability. Also, centralized training, and retraining in case of changes in the environment, require collecting local experience trajectories of each agent, i.e., observations, actions, rewards, and next observations. Transferring such information to a centralized entity to retrain these agents is inefficient compared to local retraining, especially for highly dynamic environments that keep changing over time.
@@
Determining the optimal timing for retraining the agents in a centralized entity presents a critical trade-off between decision accuracy and the significant retraining overhead. In addition, an RL decision can only be made when the observation of the current step is available, along with the reward of the previous step. Observing the state, for each decision, of every Fog node in large-scale Fog networks will cause a significant delay, especially with frequently generated workloads [4]. This delay is exponentially proportional to the number of agents in a distributed solution, making it impractical in real-world network deployment either with a centralized or a distributed solution. This delay is often not considered in the literature for Fog load-balancing solutions, assuming observations are readily available for every action at no cost. Without considering the time needed to transfer Fog load provisioning information to all RL agents, the effectiveness of these solutions in practice can not be guaranteed.
@@
Gao et al. [5] used MARL to address the scalability problem of RL-based load-balancing solutions. They provide task offloading and resource allocation in large-scale heterogeneous Multi-access Edge Computing (MEC) systems. In their design, each edge server is regarded as an actor agent while still considering a centralized critic, where the critic is a neural network used to evaluate the actions of actor networks [6]. This centralized critic facilitates the coordination of offloading policies among those distributed actors, leading to a common policy. Besides not being fully distributed, the authors assumed that observations are readily available when the agents need to make assignment decisions, where the delay of exchanging Fog status information to every agent is ignored.
@@
Baek et al. [7] proposed a centralized critic network with distributed individual actor networks. Those distributed agents are located in APs and they cooperatively learn to solve the joint link and server load-balancing problem. A common policy for the distributed cooperative agents is optimized using parameter-sharing and distributed execution; the agents communicate among themselves while evaluating a common value function. The authors assume that updated policy parameters are shared between the agents by exchanging communication messages over limited bandwidth channels. However, they did not show the transmission delay associated with those communication messages, which augments when scaling up the solution for a larger-scale Fog environment. Such a delay can have a significant impact on the collaboration of the agents due to sharing outdated parameters.
@@
Similarly, Goudarzi et al. [8] proposed a MARL framework to improve the execution delay of heterogeneous Fog services using multiple actors and a centralized critic. They deployed distributed agents (they call them brokers) that work in parallel to make service offloading decisions across different Fog nodes. A centralized service offloading model, i.e., the learner of the target policy, is trained using experience trajectories generated by the distributed actors. But again, they did not consider the delay caused by exchanging such information in a large-scale environment, especially when request rates are high. In addition, they did not consider the delay when updating the value function of each actor from the common centralized critic network.
@@
Wei et al. [9] proposed a scheme to improve vehicular Fog resource utilization in large-scale Vehicular Fog Computing (VFC) networks. They used MARL to decide on multi-tier task offloading in a distributed manner. Like most MARL-based solutions in the literature, they used the approach of centralized training with decentralized execution (CTDE). They share the agents’ parameters with a joint critic network using the Attention machine learning mechanism, leading to a joint network that incorporates the learning process of distributed gated actors that exploit local computational potentials. Likewise, Suzuki et al. [10] used CTDE to send learning samples from each agent to a centralized training entity, including task information, network link utilization, and Fog resources utilization. They unrealistically assumed all this information readily available in real-time for all agents when making offloading decisions.
@@
Zhang et al. [11] also used CTDE-based MARL to reduce the long-term average delay and energy consumption while improving the throughput of dynamic Edge networks using online Request Scheduling. Centralized training was adopted to improve the long-term system performance by realizing the implicit cooperation among edge nodes. This cooperation was obtained by transmitting the replay buffer data of each agent into a shared training server to train a common policy. During distributed execution, on the other hand, only local environment information is needed for each edge node to make the request scheduling decision. However, the authors did not discuss the scope of local information and the delay associated with transmitting such information over the network. In a large-scale network, such a delay is not tolerable for real-time decisions required by real-time applications.
@@
Lu et al. [12] proposed MARL-based dynamic scheduling to handle the massive number of service requests by the ever-increasing number of vehicles in Internet-of-Vehicle (IoV) networks. They leveraged CTDE to learn high-quality scheduling rules for the service function scheduling problem. In their architecture, the generated scheduling actions of the distributed policies are evaluated by separate critic networks deployed in a centralized controller node. Similarly, Cui et al. [13] proposed offline CTDE to cooperatively optimize the offloading strategy for computation-intensive and delay-sensitive IoV applications. Offline CTDE helped reduce the communication between the agents while increasing the stability of the environment. However, offline execution does not allow the agents to dynamically adapt to environment changes. In addition, their solution will not scale to larger networks as all their agents must observe the global state of the environment, which is assumed, unrealistically, readily available for every decision to be made.
@@
CTDE was also used by Gao et al. [14] to predict the queuing delay of edge servers by training a centralized critic in the Cloud. To evaluate each agent’s policy, the critic network needs to collect observations and their corresponding actions from all agents in the system. Then, the agents independently make their decisions using their policies distributively. Mangiaracina et al. [15] proposed adaptive MARL to manage data placement among Fog nodes considering Quality of Service (QoS) requirements. Unlike CTDE, they enabled cooperation between multiple agents using the concept of a shared repository accessible by all agents. This repository contains knowledge bases and event logs shared between those agents to improve their decision-making. The authors claim that this centralized repository does not affect the overall performance as a bottleneck in the architecture. However, they did not consider the time each agent needs to access information from this repository.
@@
To coordinate the decisions of independent agents, Yang et al. [16] proposed that the decisions of each agent must depend on shared global state information. They proposed a centralized migration scheduling model to collect server load information and calculate the migration likelihood. They estimate the global information in a centralized model to reduce the communication overhead caused by broadcasting this information from each agent to the centralized critic network. They showed a 20% gap in performance between their approach and the optimal solution but with only half the decision-making time, where the optimal solution is a centralized decision-making method having the actions of each agent and the actual state information of each Fog node (not estimated) available for all agents unrealistically in real-time.
@@
Based on the literature discussed above, most existing MARL solutions adopt a centralized training approach and rely on unrealistic assumptions regarding collecting observations (see Table I). The centralized component in these solutions raises several challenges, especially in large-scale environments, including computational complexity, exploration of larger state and action spaces, and the communication overhead due to exchanging training information with the centralized entity. A truly distributed solution remains preferable to enhance scalability and decision-making time, where agents can learn and make decisions independently without centralized or inter-agent coordination. Allowing the agents to learn independent policies based on the unique requirements of every agent can provide faster convergence compared to learning a common policy for all the agents.
@@
To fill this gap in the literature, this paper proposes a fully distributed solution using independent agents that efficiently distribute IoT load in Fog networks without reliance on centralized components or inter-agent communication. Besides mitigating communication overhead, avoiding coordination between the agents leads to learning separate decision-making policies, which is easier than learning a common load distribution strategy. These independent agents learn to reduce workload accumulation in Fog nodes by receiving a reward for every load assignment decision representing the accumulated load in the selected Fog node. Optimizing this common objective encourages each agent to leverage nearby resources before seeking remote resources, mitigating the effect of the non-stationarity problem of independently trained agents.
This reward function allows for a global-scale implementation when forming smaller groups of agents each managing a small subset of nearby Fog nodes. This grouping reduces state and action spaces for the agents, making learning easier and state observation overhead smaller. To reduce the overhead of observing Fog information further, we present the first attempt to evaluate collecting state observations and rewards using an interval-based Gossip multi-casting protocol. This approach is more realistic than assuming such information is readily available immediately before every decision. This evaluation explores the trade-off between optimality and applicability of the solution in real-world networks, leading to a deployment-ready solution.